# On the rank index of projective curves of almost minimal degree

### v1

@@ Abstract.
In this article, we investigate the rank index of projective curves $\mathcal{C}\subset\mathbb{P}^{r}$ of degree $r+1$ when $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ for the standard rational normal curve $\tilde{\mathcal{C}}\subset\mathbb{P}^{r+1}$ and a point $p\in\mathbb{P}^{r+1}\setminus\tilde{\mathcal{C}}^{3}$. Here, the rank index of a closed subscheme $X\subset\mathbb{P}^{r}$ is defined to be the least integer $k$ such that its homogeneous ideal can be generated by quadratic polynomials of rank $\leq k$. Our results show that the rank index of $\mathcal{C}$ is at most $4$, and it is exactly equal to $3$ when the projection center $p$ is a coordinate point of $\mathbb{P}^{r+1}$. We also investigate the case where $p\in\tilde{\mathcal{C}}^{3}\setminus\tilde{\mathcal{C}}^{2}$.
@@ 1. Introduction
Throughout this paper we work over an algebraically closed field $\mathbb{K}$ with $\operatorname{char}(\mathbb{K})=0$. Due to [HLMP21], a non-degenerate projective variety $X\subset\mathbb{P}^{r}$ is said to satisfy property $\operatorname{QR}(k)$ if its homogeneous ideal is generated by quadratic polynomials of rank at most $k$. When $X$ is cut out ideal-theoretically by quadrics, we define the rank index of $X$, denoted by $\operatorname{rank-index}(X)$, as the least integer $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. Thus
@@
since $X$ is irreducible and nondegenerate in $\operatorname{\mathbb{P}}^{r}$. Many classical constructions in projective geometry such as rational normal scrolls and Segre-Veronese varieties satisfy property $\operatorname{QR}(4)$ since their homogeneous ideals are generated by $2$-minors of one or several matrices of linear forms (cf. [Ber08], [H0̀2], [Puc98], [SS11], [Sul08], etc). Recently, it has been found that there are many projective varieties whose rank index is equal to $3$, which is possibly the minimal value. The $d$-th Veronese embedding of $\operatorname{\mathbb{P}}^{n}$ satisfies property $\operatorname{QR}(3)$ for all $n\geq 1$ and $d\geq 2$ if $\operatorname{char}(\mathbb{K})\neq 3$ (cf. [HLMP21]). Every linearly normal curve of arithmetic genus $g$ and degree $\geq 4g+4$ satisfies property $\operatorname{QR}(3)$ (cf. [Par22]). Also, the authors in [MP23] found the rank index of various projective varieties, including rational normal scrolls and del Pezzo varieties. For example, if $X$ is a smooth rational normal scroll of dimension $\geq 2$, then $\operatorname{rank-index}(X)=4$.
The main purpose of this paper is to investigate the rank index of projective curves $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ of degree $r+1$. To put things in perspective, we recall some of the previously known results regarding this problem. To give precise statements, we require some notation and definitions. Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be a nondegenerate projective curve of degree $d$. Then $d\geq r$ and equality is attained if and only if it is a rational normal curve, which is obviously very well-understood. In particular, it holds that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [HLMP21, Corollary 2.4]). The next case, i.e., $d=r+1$, is exactly one of the following two cases:
  1. Case $1$ $\mathcal{C}$ is a linearly normal curve of arithmetic genus one;
  2. Case $2$ $\mathcal{C}$ is an isomorphic projection of a rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$.
In the former case, it is already known that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [Par22, Theorem 1.1]). In the latter case, let
@@ 1. Case $1$ $\mathcal{C}$ is a linearly normal curve of arithmetic genus one;
be the standard rational normal curve of degree $r+1$, and let $\tilde{\mathcal{C}}^{k}$ denote the $k$-th join of $\tilde{\mathcal{C}}$ with itself. That is, $\tilde{\mathcal{C}}^{k}$ is the closure of the set of points lying in $(k-1)$-dimensional linear subspaces spanned by general collections of $k$ points in $\tilde{\mathcal{C}}$. For joins of projective varieties, we refer to [Zak04]. Then
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$
@@
$\mbox{rank}_{\tilde{\mathcal{C}}}(p)=\min\{~{}k~{}\mid~{}p\in\tilde{\mathcal{C}}^{k}~{}\},$
is called the rank of $p$ with respect to $\tilde{\mathcal{C}}$, or simply $\tilde{\mathcal{C}}$-rank of $p$. Note that if $p_{c}$ is the $c$-th coordinate point of $\mathbb{P}^{r+1}$ for some $0\leq c\leq\lfloor(r+1)/2\rfloor$, then $\mbox{rank}_{\tilde{\mathcal{C}}}(p)=c+1$ (cf. [Par07, Section 5]). It is well known that various properties of $\mathcal{C}$ are determined by the rank of $p$ with respect to $\tilde{\mathcal{C}}$. For example, if $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{C}$, then $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ is an isomorphic projection of $\widetilde{\mathcal{C}}$ if and only if $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 3$. Moreover, the graded Betti numbers of $\mathcal{C}$ are uniquely determined by $\mbox{rank}_{\tilde{\mathcal{C}}}(p)$ (cf. [Par07] and [LP16]). In particular, the homogeneous ideal of $\mathcal{C}$ is generated by quadratic polynomials if and only if $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$. Also there is a surprising connection between some invariants of $\mathcal{C}$ in real algebraic geometry and the rank of $p$ with respect to $\tilde{\mathcal{C}}$ (cf. [BCJ24]).
@@
For the standard rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$ and a point $p\in\operatorname{\mathbb{P}}^{r+1}$ with $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$, let
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})\subset\operatorname{\mathbb{P}}^{r}$
@@
    Suppose that $p=p_{c}$ is the $c$-th coordinate point of $\mathbb{P}^{r+1}$ (and hence $23\leq c\leq r-$ 2). Then $\operatorname{rank-index}(\mathcal{C})=3$.
The proofs of Theorem 1.1. (1) and (2) are respectively in Section 3 and Section 4.
One of the reasons why the problem of determining the rank index of the curve $\mathcal{C}$ in Theorem 1.1 is interesting is that, unlike previously known varieties satisfying property $\operatorname{QR}(3)$, which are all linearly normal, there is the difficulty that $\mathcal{C}$ is not linearly normal and hence its properties change continuously as the projection center $p$ varies.
In the proof of Theorem 1.1.(1), we use the fact that there is a unique rational normal surface scroll $\mathcal{S}$ that contains $\mathcal{C}$. Indeed, the homogeneous ideal of $\mathcal{C}$ is generated by that of $\mathcal{S}$ and the quadratic polynomials obtained from the equation of $\mathcal{C}$ as a divisor of $\mathcal{S}$. Note that the existence of such a surface $\mathcal{S}$ can be found in [EEK88, Proposition 1.2] and [BS07, Theorem 7.3] and the uniqueness of $\mathcal{S}$ is recently proven in [KMP24, Theorem 5.2]. We provide a new constructive proof of the existence of $\mathcal{S}$ by using the apolar ideal of the binary form corresponding to the projection center $p$ in Section 3.
The proof of Theorem 1.1. (2) is obtained by combining the framework of induction on the degree of $\mathcal{C}$ with the investigation of rank $3$ quadratic equations of $\tilde{\mathcal{C}}$ that are also contained in the homogeneous ideal of $\mathcal{C}$.
Regarding Theorem 1.1, one can naturally ask whether there exists a point $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\tilde{\mathcal{C}}^{3}$ such that the rank index of $\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})$ is equal to $4$. We do not have such an example. On the contrary, Theorem 1.1.(2) leads us to formulate the following conjecture.
@@
Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be as in Theorem 1.1. Then $\operatorname{rank-index}(\mathcal{C})=3$.
@@
is generated by the quadratic polynomials in the homogeneous ideal of $\mathcal{C}$. Note that $\mathcal{C}$ is $3$-regular and hence $\mbox{length}(\mathcal{C}\cap\mathbb{L})=3$ (cf. [HSV91]). Regarding the rank index of $X$, we obtain the following interesting partial result.
@@
Suppose that $r\geq 4$. Let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as above. Then
@@
The proofs of Theorem 1.3 is in Section 5.
By Theorem 1.3, it remains to determine the rank index of $X$ when $\mathcal{C}\cap\mathbb{L}$ consists of three simple points. There are many example such that $X$ can satisfy property $\operatorname{QR}(3)$ and hence $\operatorname{rank-index}(X)=3$. This leads us to formulate the following conjecture.
@@
Let $\mathcal{C}$, $\mathbb{L}$ and $X=\mathcal{C}\cup\mathbb{L}$ be as in Theorem 1.3. Then $\operatorname{rank-index}(X)=3$ if and only if $\mathcal{C}\cap\mathbb{L}$ consists of three simple points.
Obviously, Theorem 1.3 proves one direction of 1.4.
@@
Section 2 outlines key definitions and foundational concepts including (smooth) curves of almost minimal degree, apolarity of binary forms, and the rank index of varieties. Section 3 presents our primary results on the rank index for projected curves. In Section 4, we focus on the quadratic rank index of projected curves obtained by linear projection from a standard basis in projective spaces. Section 5 includes explicit calculations for certain projective schemes not covered in the previous sections that provide additional supports for conjectures concerning the rank index.
Acknowledgement. The first named author was supported by the Institute for Basic Science (IBS-R032-D1-2023-a00). The second named author was supported by a KIAS Individual Grant(MG083101) at Korea Institute for Advanced Study. The third named author was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2022R1A2C1002784).
@@ 2. Preliminaries
2.1. Apolarity on binary forms
Given a polynomial ring $\mathbb{K}[s,t]$ with the standard grading i.e., $\deg(s)=\deg(t)=1$. A monomial is denoted by $s^{a}t^{b}$ for nonnegative integers $a$ and $b$. We define the degree of a monomial $s^{a}t^{b}$ as $a+b$.
@@ 2.1. Apolarity on binary forms
Let $R=\mathbb{K}[s,t]$ be the graded ring over $\mathbb{K}$ with the standard grading on the variables. That is, $R=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}R_{i}$ where $R_{i}$ is the space of homogeneous polynomials of degree $i$. Let $\mathcal{R}=\mathbb{K}[S,T]$ be a graded ring whose graded components $\mathcal{R}_{i}$ are the $\mathbb{K}$-vector spaces of homogeneous polynomials of degree $i$.
We consider a bilinear map of vector spaces
$\mathcal{R}_{j}\times R_{i}\rightarrow R_{i-j}$
defined by
$(F,f)\mapsto F\circ f=\sum_{0\leq k\leq j}c_{k}\frac{\partial^{j}}{\partial s^{j-k}\partial t^{k}}f$
where $F=\sum_{0\leq k\leq j}c_{k}S^{j-k}T^{k}$. The apolar ideal of a homogeneous polynomial $f\in R$ is defined as the set of homogeneous polynomials $F\in\mathcal{R}$ such that $F\circ f=0$, denoted by $(f)^{\perp}$. In other words,
$(f)^{\perp}:=\{F\in\mathcal{R}:F\circ f=0\}.$
@@
(cf. [IK99, Theorem 1.44.]) Suppose $f\in\mathbb{K}[s,t]$ is a binary form of degree $d$. Then, $(f)^{\perp}$ is a complete intersection ideal generated by two binary forms in $\mathcal{R}$. This means that if we let $F_{l}$ be the lower-degree generator and $F_{h}$ the higher-degree generator of $(f)^{\perp}\subset\mathbb{K}[S,T]$, then
@@
where $V(F_{l},F_{h})=\emptyset$ and $\deg F_{l}+\deg F_{h}=d+2$ with $\deg F_{l}\leq\deg F_{h}$. If $\deg F_{l}<\deg F_{h}$, then $F_{l}$ is uniquely determined by $f$, while $F_{h}$ is unique modulo the principal ideal $\langle F_{l}\rangle$.
For example, if $f=\alpha s+\beta t\in\mathbb{K}[s,t]$ is a linear binary form, then the lower-degree generator of $(f)^{\perp}$ is $F_{l}=\beta S-\alpha T\in\mathcal{R}$ and $F_{h}$ is any quadratic forms in $\mathcal{R}$.
The apolarity of forms is a useful tool for studying the decomposition of forms into sums of powers of linear forms. This decomposition is known as the Waring decomposition or the (generalized) additive decomposition of forms. For more details on the generalized additive decomposition and apolarity, see [IK99].
We remark that given a vector space $R_{d}$ for a fixed integer $d$, the pairing $R_{d}\times\mathcal{R}_{d}\rightarrow R_{0}=\mathbb{K}$ forms an inner product on these vector spaces. Therefore, we can identify $\mathcal{R}_{d}$ with the dual vector space of $R_{d}$ via the differential operator. In particular, for $f\in R_{d}$ and $g\in\mathcal{R}_{d}$, the operation $g\circ f=f\circ g$ is defined by
@@
where $\delta_{ij}$ is the Kronecker delta. Therefore, any form $f(s,t)$ can be expressed as

### v2

@@ Abstract.
[…]
@@
Corresponding author: Jaewoo Jung (jaewoojung@dgist.ac.kr)
@@ 1. Introduction
Throughout this paper we work over an algebraically closed field $\mathbb{K}$ with $\operatorname{char}(\mathbb{K})=0$. According to [8], a non-degenerate projective variety $X\subset\mathbb{P}^{r}$ is said to satisfy property $\operatorname{QR}(k)$ if its homogeneous ideal is generated by quadratic polynomials of rank at most $k$. When $X$ is cut out ideal-theoretically by quadrics, we define the rank index of $X$, denoted by $\operatorname{rank-index}(X)$, as the least integer $k$ such that $X$ satisfies property $\operatorname{QR}(k)$. Thus
@@
since $X$ is irreducible and nondegenerate in $\operatorname{\mathbb{P}}^{r}$. Many classical constructions in projective geometry such as rational normal scrolls and Segre-Veronese varieties satisfy property $\operatorname{QR}(4)$ since their homogeneous ideals are generated by $2$-minors of one or several matrices of linear forms (cf. [1], [7], [17], [18], [19], etc). Recently, it has been found that there are many projective varieties whose rank index is equal to $3$, which is the minimal value for irreducible varieties. The $d$-th Veronese embedding of $\operatorname{\mathbb{P}}^{n}$ satisfies property $\operatorname{QR}(3)$ for all $n\geq 1$ and $d\geq 2$ if $\operatorname{char}(\mathbb{K})\neq 3$ (cf. [8]). Every linearly normal curve of arithmetic genus $g$ and degree $\geq 4g+4$ satisfies property $\operatorname{QR}(3)$ (cf. [16]). Also in [13] the authors found the rank index of various projective varieties, including rational normal scrolls and del Pezzo varieties. For example, if $X$ is a smooth rational normal scroll of dimension $\geq 2$, then $\operatorname{rank-index}(X)=4$.
The main purpose of this paper is to investigate the rank index of projective curves $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ of degree $r+1$. To put things in perspective, we recall some of the previously known results regarding this problem. To give precise statements, we require some notation and definitions. Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be a nondegenerate projective curve of degree $d$. Then $d\geq r$ and equality is attained if and only if $\mathcal{C}$ is a rational normal curve, which is obviously very well-understood. In particular, it holds that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [8, Corollary 2.4]). If $d=r+1$, then the following two cases may occur:
  1. Case $1.$ $\mathcal{C}$ is a linearly normal curve of arithmetic genus one;
  2. Case $2.$ $\mathcal{C}$ is an isomorphic projection of a rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$.
In the former case, it is already known that $\operatorname{rank-index}(\mathcal{C})=3$ (cf. [16, Theorem 1.1]). In the latter case, let
@@
be the standard rational normal curve of degree $r+1$, and let $\tilde{\mathcal{C}}^{k}$ denote the $k$-th self-join of $\tilde{\mathcal{C}}$. That is, $\tilde{\mathcal{C}}^{k}$ is the closure of the set of points lying in (k1)(k-1)-dimensional linear subspaces spanned by general collections of $k$ points in $\tilde{\mathcal{C}}$. For joins of projective varieties, we refer to [20]. Then our rational curve $\mathcal{C}\subseteq\mathbb{P}^{r}$ of degree $d=r+1$ may be written in the form:
[…]
@@
$\mbox{rank}_{\tilde{\mathcal{C}}}(p)=\min\{k\mid p\in\tilde{\mathcal{C}}^{k}\},$
[…]
@@
For the standard rational normal curve $\widetilde{\mathcal{C}}\subset\operatorname{\mathbb{P}}^{r+1}$ and a point $p\in\operatorname{\mathbb{P}}^{r+1}\setminus\widetilde{\mathcal{C}}^{2}$ with $\mbox{rank}_{\tilde{\mathcal{C}}}(p)\geq 4$, let
$\mathcal{C}=\pi_{p}(\tilde{\mathcal{C}})\subseteq\operatorname{\mathbb{P}}^{r}$
@@
    Suppose that $p=p_{c}$ is the c-th coordinate point of $\mathbb{P}^{r+1}$ (and hence $3\leq c\leq r-2$). Then $\operatorname{rank-index}(\mathcal{C})=3$.
The proofs of statements (1) and (2) of Theorem˜1.1 are given respectively in Section˜3 and Section˜4.
One of the main reasons why determining the rank index of the curve $\mathcal{C}$ in Theorem˜1.1 is interesting is that, in contrast to all previously studied varieties satisfying property $\operatorname{QR}(3)$, which are linearly normal, the curve $\mathcal{C}$ is not linearly normal. As a consequence, its geometric and algebraic properties depend continuously on the choice of the projection center $p$.
In the proof of Theorem˜1.1(1), we use the fact that there is a unique rational normal surface scroll $\mathcal{S}$ that contains $\mathcal{C}$. Indeed, the homogeneous ideal of $\mathcal{C}$ is generated by that of $\mathcal{S}$ and the quadratic polynomials obtained from the equation of $\mathcal{C}$ as a divisor of $\mathcal{S}$. Note that the existence of such a surface $\mathcal{S}$ can be found in [4, Proposition 1.2] and [3, Theorem 7.3] and the uniqueness of $\mathcal{S}$ is recently proven in [11, Theorem 5.2]. We provide a new constructive proof of the existence of $\mathcal{S}$ by using the apolar ideal of the binary form corresponding to the projection center $p$ in Section˜3.
The proof of Theorem˜1.1(2) is obtained by combining induction on the degree of $\mathcal{C}$ with the investigation of rank $3$ quadratic equations of $\tilde{\mathcal{C}}$ that are also contained in the homogeneous ideal of $\mathcal{C}$.
[…]
@@
Let $\mathcal{C}\subset\operatorname{\mathbb{P}}^{r}$ be as in Theorem˜1.1. Then $\operatorname{rank-index}(\mathcal{C})=3$.
@@
is generated by the quadratic polynomials in the homogeneous ideal of $\mathcal{C}$. (See Lemma˜5.1 for details.) Note that $\mathcal{C}$ is $3$-regular (cf. [9, Theorem 3]). Regarding the rank index of $X$, we obtain the following interesting partial result.
@@
Suppose that $r\geq 4$. Assume that $\operatorname{rank}_{\widetilde{\mathcal{C}}}(p)=3$ and let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as above. Then
@@
The proof of Theorem˜1.3 is given in Section˜5.
By Theorem˜1.3, it remains to determine the rank index of $X$ when $\mathcal{C}\cap\mathbb{L}$ consists of three simple points. There are many example such that $X$ has property $\operatorname{QR}(3)$ and hence satisfies $\operatorname{rank-index}(X)=3$. This leads us to the following conjecture.
@@
Let $\mathcal{C}$, $\mathbb{L}$, and $X=\mathcal{C}\cup\mathbb{L}$ be as in Theorem˜1.3. Then $\operatorname{rank-index}(X)=3$ if and only if $\mathcal{C}\cap\mathbb{L}$ consists of three distinct reduced points.
Obviously, Theorem˜1.3 proves one direction of ˜1.4.
@@
Section˜2 outlines key definitions and foundational concepts including (smooth) curves of almost minimal degree, apolarity of binary forms, and the rank index of varieties. Section˜3 presents our primary results on the rank index for projected curves. In Section˜4, we focus on the quadratic rank index of projected curves obtained by linear projection from a standard basis in projective spaces. Section˜5 includes explicit calculations for certain projective schemes not covered in the previous sections that provide additional support for conjectures concerning the rank index.
@@ 2. Preliminaries
2.1. Apolarity of binary forms
In this subsection we provide some basics related to the notion of apolarity of binary forms, which will play an essential role later. To do so, we fix some notation which will be kept throughout: Let $R:=\mathbb{K}[s,t]$ and $\mathcal{R}:=\mathbb{K}[S,T]$ be two bivariate standard graded polynomial rings over the field $\mathbb{K}$, so that $\deg(s)=\deg(t)=\deg(S)=\deg(T)=1$, $R=\bigoplus_{i\in\mathbb{Z}}R_{i},\mathcal{R}=\bigoplus_{i\in\mathbb{Z}}\mathcal{R}_{i}$ where $R_{i}:=\bigoplus_{j=0}^{i}\mathbb{K}s^{j}t^{i-j}$ and $\mathcal{R}_{i}:=\bigoplus_{j=0}^{i}\mathbb{K}S^{j}T^{i-j}$ are the $i$-th graded components of $R$ respectively of $\mathcal{R}$ for all $i\in\mathbb{Z}$. As usual, we call the polynomials $f\in R_{i}$ and $F\in\mathcal{R}_{i}$ binary forms of degree $i$ in $R$ respectively in $\mathcal{R}$. Correspondingly, we respectively call the sets $R^{\mathrm{hom}}:=\bigcup_{i\in\mathbb{Z}}R_{i}$ and $\mathcal{R}^{\mathrm{hom}}:=\bigcup_{i\in\mathbb{Z}}\mathcal{R}_{i}$ of homogeneous polynomials in $R$ and in $\mathcal{R}$ the sets of binary forms (in $R$ and in $\mathcal{R}$).
@@ 2.1. Apolarity of binary forms
For all $i,j\in\mathbb{Z}$ we consider the bilinear map of $\mathbb{K}$-vector spaces
$\circ_{(j,i)}:\mathcal{R}_{j}\times R_{i}\to R_{i-j}\quad\text{and the induced map}\quad\circ:\mathcal{R}^{\mathrm{hom}}\times R^{\mathrm{hom}}\to R^{\mathrm{hom}}$
where $\circ_{(j,i)}$ is defined by
$(F,f)\mapsto F\circ_{(j,i)}f=F\circ f:=\sum_{k=0}^{j}c_{k}\frac{\partial^{j}}{\partial s^{j-k}\partial t^{k}}f$
where $F=\sum_{k=0}^{j}c_{k}S^{j-k}T^{k}$. The apolar ideal of a homogeneous polynomial $f\in R$ is defined to be the ideal generated by all homogeneous polynomials $F\in\mathcal{R}$ for which $F\circ f=0$, and we denote it by $(f)^{\perp}$. In other words,
$(f)^{\perp}:=(F\in\mathcal{R}:F\circ f=0).$
@@
(cf. [10, Theorem 1.44.]) Suppose $f\in\mathbb{K}[s,t]$ is a binary form of degree $d$. Then, $(f)^{\perp}$ is a complete intersection ideal generated by two binary forms in $\mathcal{R}$. More precisely: If $F_{\ell},F_{h}$ are of minimal respectively maximal degree among all $F\in\mathcal{R}_{\mathrm{hom}}\setminus\{0\}$ with $F\circ f=0$, then
@@
where $V(F_{l},F_{h})=\emptyset$ and $\deg F_{l}+\deg F_{h}=d+2$ with $\deg F_{l}\leq\deg F_{h}$. If $\deg F_{l}<\deg F_{h}$, then $F_{\ell}$ is essentially uniquely determined by $f$, while $F_{h}$ is essentially unique modulo the principal ideal $\langle F_{l}\rangle$.
For example, if $f=\alpha s+\beta t\in\mathbb{K}[s,t]$ is a linear binary form, then the lower-degree generator of $(f)^{\perp}$ is $F_{l}=\beta S-\alpha T\in\mathcal{R}$ and $F_{h}$ is any quadratic form in $\mathcal{R}$.
The apolarity of forms is a useful tool for studying the decomposition of forms into sums of powers of linear forms. This decomposition is known as the Waring decomposition or the (generalized) additive decomposition of forms. For more details on the generalized additive decomposition and apolarity, see [10].
We remark that for any integer $d\geq 0$, the pairing $\circ:\mathcal{R}_{d}\times R_{d}\to R_{0}=\mathbb{K}$ is perfect and hence forms a non-degenerate inner product of the two vector spaces $\mathcal{R}_{d}$ and $R_{d}$, meaning that for any basis $v_{0},v_{1},\ldots,v_{d}$ of $\mathcal{R}_{d}$ and any basis $w_{0},w_{1},\ldots,w_{d}$ of $R_{d}$ the matrix $v_{\bullet}\circ w_{\bullet}:=(v_{i}\circ w_{j}\mid 0\leq i,j\leq d)\in\mathbb{K}^{(d+1)\times(d+1)}$ is of rank $d+1$. Therefore, we can identify $\mathcal{R}_{d}$ with the dual vector space of $R_{d}$ via the differential operator. In particular, for $f\in R_{d}$ and $g\in\mathcal{R}_{d}$, the inner product $g\circ f$ is given by
@@
where $\delta_{ij}$ is the Kronecker delta. Therefore, any form $f(s,t)\in\mathbb{K}[s,t]_{d}$ can be expressed as